OpenStudy (anonymous):

Can someone please help me? :) Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (anonymous):

ok so i am trying to find some site that will help me but i only in 7th but now a little 10th grade math

OpenStudy (anonymous):

Oh ok lol

OpenStudy (anonymous):

Find anything? lol :)

OpenStudy (anonymous):

no but i did find how to do it on microsoft

OpenStudy (anonymous):

so x -2 = 2-2=0 because (2,5) then 3-2 =1

OpenStudy (anonymous):

but i not sure on the y axis

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem